# A General Framework for Joint Multi-State Models

### v3

@@ 7.2.1 Inference
Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing. The model was fitted using the Adam optimizer. The optimization process was run until convergence based on the same criterion as the simulation study with a tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
Before displaying the fitted parameters, the Fisher Information Matrix was computed, implementing the method described in [baey2023efficient], directly taking into account shared parameters. The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix. The results are summarized in Table 8.
@@
$\gamma_{1}$ 2.1600 0.0470 $\gamma_{2}$ 0.1590 0.0140
$\gamma_{3}$ 1.0010 0.0030 $\tilde{Q}_{1}$ -0.5700 0.0430
$\tilde{Q}_{2}$ 0.9570 0.0960 $\tilde{Q}_{3}$ 0.3820 0.0430
$\tilde{Q}_{4}$ -0.2050 0.1570 $\tilde{Q}_{5}$ -2.3310 0.1380
$\tilde{Q}_{6}$ 0.5220 0.0470 $\tilde{R}$ 2.0830 0.0260
$\alpha^{3\mid 1}$ 0.7220 0.7150 $\alpha^{1\mid 0}$ -0.5400 0.4030
$\alpha^{2\mid 1}$ -1.0030 0.2000 $\alpha^{2\mid 0}$ -0.1210 0.0380
$\alpha^{3\mid 2}$ -0.0480 0.2520 $\alpha^{3\mid 0}$ -0.3070 0.0740
$\beta_{1}$ -0.2990 0.9130 $\beta_{2}$ 0.0970 1.0300
@@
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (variance parameters excluded).
Notably, the estimated standard errors are consistent with the number of observations per transition, with larger standard errors for transitions with fewer observations. Therefore, parameters with too few observations should be interpreted with caution, if at all. The linear covariate parameters also exhibit large standard errors, which indicates the coefficients are not significative in this setting.
The association parameters $\alpha$ being negative, we can infer that cognitive decline (i.e. rate of change) is associated with an increase in physical dependency.
7.2.2 Dynamic Prediction
Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction. For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation. Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 9.
Formally, the prediction function $\chi_{u}^{*}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as
$\text{accuracy}(u)=\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\chi_{u\wedge C_{i}}(\mathcal{T}_{i})=s_{i}}.$
Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\chi_{u\wedge C_{i}}(\mathcal{T}_{i})$ may be computed based on the censored trajectories. However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time. Nevertheless, this inclusion is necessary to ensure fairness in accuracy metrics, since individuals who are not censored are generally more likely to have progressed to more advanced states of physical dependency than those who are censored.
Figure 9: Accuracy of the predicted state at each time point with respect to the true state for different truncation times.
Here, the initial drops correspond to the truncation times, before which the accuracy is always $100\%$.
As expected from dynamic predictions, the accuracy increases as more data is available, therefore when the truncation time grows. We also observe that the longer the prediction horizon, the lower the accuracy of the predictions.
@@ 8 Conclusion
Since the model may potentially involve many transitions and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters. This is especially beneficial in settings with limited data for certain transitions, or when biological or clinical knowledge suggests similar effects across multiple transitions. Parameter sharing improves statistical efficiency, reduces overfitting, and facilitates model interpretability.
@@ Appendix A Multi-State Simulation
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property by considering one transition at a time, according to Algorithm 2. The algorithme is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation.
@@
1:subject-specific censoring time $C_{i}\in\mathbb{R}\cup\{+\infty\}$; covariates $X_{i}$; latent variables $b_{i}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i}^{\text{surv}}$, defaults to $t_{i}^{\text{surv}}=$-$\infty$.
@@
3:Initialize $\ell_{i}\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$.
@@
5:while $T_{i(\ell_{i}-1)}<C_{i}$ and $\{s:(S_{i(\ell_{i}-1)},s)\in E\}\neq\emptyset$ do
6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
7:   Draw $T_{i\ell_{i}}^{s}$ such that
8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid b_{i},X_{i},T_{i(\ell_{i}-1)},\theta)\,dw,$
9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
10:  end for
11:
12:  Set $T_{i\ell_{i}}\leftarrow\min_{m}T_{i\ell_{i}}^{s}$, and $S_{i\ell_{i}}\leftarrow\arg\min_{m}T_{i\ell_{i}}^{s}$.
13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
14:  $\ell_{i}\leftarrow\ell_{i}+1$.
15:end while
16:
17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
19:end if
20:return $\mathcal{T}_{i}$
@@
First consider the case $t_{i}^{\text{surv}}=$-$\infty$. Let $t>T_{i\ell_{i}}$:
$\mathbb{P}(T_{i(\ell_{i}+1)}>t\mid T_{i\ell_{i}},S_{i\ell_{i}})$ $=\mathbb{P}(\bigcap_{s:\,(E_{i(\ell_{i}-1)},s)\in E}T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\prod_{s:\,(S_{i\ell_{i}},s)\in E}\mathbb{P}(T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
Furthermore:
$\mathbb{P}(S_{i(\ell_{i}+1)}\mid T_{i(\ell_{i}+1)}=t,T_{i\ell_{i}},S_{i\ell_{i}})$
$\propto p(\{T_{i(\ell_{i}+1)}^{S_{i(\ell_{i}+1)}}=t\}\cap\bigcap_{\begin{subarray}{c}s:\,(S_{i\ell_{i}},s)\in E,\\ s\neq S_{i(\ell_{i}+1)}\end{subarray}}\{T_{i(\ell_{i}+1)}^{s}\geq t\}\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
Combining both, we retrieve the same joint density.
If $t_{i}^{\text{surv}}>$-$\infty$, one can also check that both densities are equal. Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
@@ Appendix B Expression of the Semi-Markov likelihood
The proof is short and relies on the fact that, using the semi-Markov property 2
$p\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i},\theta\right)$ $=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell},\theta\right),$
$=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell}),\theta\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise
$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right),$
and so we recover the formula cited above with the convention that an empty sum is 0.
Appendix C Stochastic optimization process
(a) $\gamma$
(b) $Q$
(c) $R$
(d) $\alpha^{1\mid 0}$
(e) $\alpha^{2\mid 0}$
(f) $\alpha^{2\mid 1}$
(g) $\beta^{1\mid 0}$
(h) $\beta^{2\mid 0}$
(i) $\beta^{2\mid 1}$
Figure 10: Evolution of the parameters during the optimization of the marginal log-likelihood using stochastic gradient descent. The dotted lines correspond to the true values.

### v4

@@ 7.3 Results
In this study, the full dataset was used to estimate the inferential results presented in Subsection 7.3.1. For the dynamic prediction analysis in Subsection 7.3.2, we employed a 5-fold cross-validation procedure, corresponding to splits of $400$ observations for model training and $100$ for validation in each fold, as to obtain unbiased estimates of the predictive performance metrics.
@@ 7.3.1 Inference
The model was fitted on the complete dataset using the Adam optimizer and the optimization process was run until convergence based on the Equation 6.2 as in the simulation study. The Fisher Information Matrix was computed, implementing the method described in [baey2023efficient]. The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix. Since, under the chosen parametrization, the value 0 lies in the interior of the admissible parameter space, Wald statistics can be used to test the nullity of each coefficient through the usual asymptotic normal approximation. However, we report only the significance results for the linear coefficients coefficients in either the covariates or the link functions, as they allow a direct interpretation. They are indicated in bold in the summarized Table 8. The optimization process itself is provided in the Appendix (see Figure 11).
@@
$\gamma_{1}$ 1.201 0.035 $\boldsymbol{\gamma_{2}}$ 0.221 0.042
$\boldsymbol{\gamma_{3}}$ 0.155 0.041 $\gamma_{4}$ 3.321 0.076
$\gamma_{5}$ 0.181 0.094 $\boldsymbol{\gamma_{6}}$ 0.568 0.119
$\gamma_{7}$ 0.636 0.188 $\boldsymbol{\gamma_{8}}$ 0.663 0.229
$\boldsymbol{\gamma_{9}}$ 1.168 0.233 $\tilde{Q}_{1}$ 2.537 0.079
$\tilde{Q}_{2}$ -0.032 0.321 $\tilde{Q}_{3}$ 0.644 0.086
$\tilde{Q}_{4}$ -1.909 0.134 $\tilde{Q}_{5}$ -0.752 0.100
$\tilde{Q}_{6}$ -0.690 0.065 $\frac{1}{2}\log(R)$ 3.023 0.008
$\log\lambda^{0\to 1}$ -1.174 1.090 $\log k^{0\to 1}$ 0.269 0.125
$\log\lambda^{0\to 2}$ 1.160 0.585 $\log k^{0\to 2}$ 0.648 0.142
$\log\lambda^{0\to 3}$ 1.663 0.769 $\log k^{0\to 3}$ 0.876 0.352
$\log\lambda^{1\to 2}$ 1.260 0.519 $\log k^{1\to 2}$ 0.434 0.097
$\log\lambda^{1\to 3}$ 2.744 0.926 $\log k^{1\to 3}$ 0.520 0.283
$\log\lambda^{2\to 3}$ 2.239 0.346 $\log k^{2\to 3}$ 0.509 0.107
$\alpha^{0\to 1}$ 1.254 1.614 $\boldsymbol{\alpha^{0\to 2}}$ -3.844 1.281
$\boldsymbol{\alpha^{0\to 3}}$ -8.487 1.362 $\boldsymbol{\alpha^{1\to 2}}$ -2.167 0.898
$\boldsymbol{\alpha^{1\to 3}}$ -8.393 0.957 $\boldsymbol{\alpha^{2\to 3}}$ -5.315 0.516
$\beta_{1}$ -0.159 0.315 $\beta_{2}$ 0.152 0.233
@@
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (only the linear coefficients included).
All link parameters $\alpha^{k\to k^{\prime}}$ for transitions $(k,k^{\prime})\in E$ are negative and statistically significant, except that from state 0 to $1$, and the most negative coefficients correspond to transitions reflecting rapid deterioration ($0\to 3$, $1\to 3$, and $2\to 3$). This pattern indicates that the current value of the MMSE score strongly influences the risk of progressing toward physical dependence, and lower cognitive performance is associated with a substantially increased transition intensity, especially for direct or fast declines.
Consistent with other analyses conducted on this dataset, both education level (CEP) and female gender are associated with more favorable cognitive and functional trajectories. Specifically, the inferred values of $\gamma_{2}$, $\gamma_{5}$, and $\gamma_{8}$ indicate that individuals with a primary school certificate (CEP) not only start from a significantly higher baseline level of MMSE, but also experience a delayed and slower progression of dementia. Similarly, the positive values of $\gamma_{3}$, $\gamma_{6}$, and $\gamma_{9}$ suggest the same remarks for women relative to men. Regarding the association coefficient $\beta$, the relatively large standard errors imply that the remaining effects are modest and that the link function already captures most of the relevant information.
7.3.2 Multi-State Dynamic Prediction
After fitting the model on each training fold, we performed multi-state dynamic prediction on the corresponding validation fold. For each validation fold $V_{k}$, with $k\in\llbracket 5\rrbracket$, predictions were computed using a single Monte Carlo integration scheme as described in Section 5. We assessed the performance of our model using two metrics: the Brier Skill Score (BSS), which is one of the most standard metric in this context, as well as the accuracy. More specifically, for each new individual $i$, longitudinal measurements and trajectories were truncated at various landmarks $t$, and predictions were made for future time points $u\geq t$ beyond each truncation. We considered the quantity $\chi_{u}^{*}(\mathcal{T}_{i}^{*})$ as defined by Example 1, which is estimated using the information available up to $t$, i.e., $\left(\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right)$.
Because the complete trajectories are not fully observed in the test sample and the quantity $\chi_{u}^{*}(\mathcal{T}_{i}^{*})$ is not available for $u>C_{i}$, predictions must account for censoring. For a given landmark time $t$, we obtain a collection of probability estimates
$\{\hat{\chi}_{i,u}(t)\}\subset\Delta\left(\{0,1,2,3\}\right),$
each computed from the information available up to $t$. The Brier Skill Score for the fold $V_{k}$ with $k\in\llbracket 5\rrbracket$ is then defined by
$\mathrm{BSS}_{k}(t,u)\coloneqq 1-\frac{\mathrm{BS}_{k}(t,u)}{\mathrm{BS}_{k}^{(0)}(u)},$
where
$\begin{cases}\mathrm{BS}_{k}(t,u)\coloneqq\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\left\|\hat{\chi}_{i,u\wedge C_{i}}(t)-\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right\|^{2},\\ \mathrm{BS}_{k}^{(0)}(u)\coloneqq\min_{p}\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\left\|p-\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right\|^{2}.\end{cases}$
Hence, values of $\mathrm{BSS}_{k}(t,u)$ close to $1$ indicate a substantial improvement over the optimal constant prediction.
Similarly, the accuracy at $(t,u)$ on the fold $V_{k}$ is defined as
$\mathrm{Acc}_{k}(t,u)\coloneqq\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\mathds{1}\left\{\operatorname*{\arg\max}_{j\in\{0,1,2,3\}}\left(\hat{\chi}_{i,u\wedge C_{i}}(t)\right)_{j}=\operatorname*{\arg\max}_{j\in\{0,1,2,3\}}\left(\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right)_{j}\right\},$
which corresponds to the proportion of correctly identified states.
Figure 9 illustrates our dynamic predictions framework evaluated at three distinct landmark times using both metrics estimated from our cross-validation procedure. Both metrics are averaged across the $5$ folds.
Figure 9: Average values on the 55-fold cross-validation of the Brier Skill Score (top) and accuracy (bottom) at each time point for different prediction landmarks (dotted lines). Colored bands correspond to $\mathrm{mean}\pm 1.65\mathrm{sd}$.
Across the three prediction landmarks, the Brier Skill Score remains above $40\%$ at $10$ years beyond the prediction landmark. This indicates a very strong improvement over the best constant baseline model, despite the limited information available at the time of prediction.
Similarly, the classification accuracy consistently exceeds $60\%$, often higher, thus the model correctly identifies the majority of states. This level of performance persists even for long-term horizons extending more than $20$ years into the future, including during the early phases of the study.
As expected, predictions made immediately following each truncation are the most precise, while uncertainty naturally increases as the prediction horizon extends further. Both metrics also get substantially better as the prediction landmarks include more data. Both the Brier Skill Score (BSS) and classification accuracy remain robust over time, providing strong evidence of the reliability and stability of our predictive framework.
@@ 8 Conclusion
Since the model may potentially involve many transitions, and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters. This is especially beneficial in settings with limited data for certain transitions, or when biological or clinical knowledge suggests similar effects across multiple transitions. Parameter sharing improves statistical efficiency, reduces overfitting, and facilitates model interpretability.
@@ Appendix A Multi-State Simulation
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property 1 by considering one transition at a time, according to Algorithm 3. The algorithm is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i,\mathrm{cond}}$ that uses the Chasles relation.
Here, we deliberately avoid imposing any conditioning in order to preserve maximal generality, since the algorithm is intended to be applicable beyond the specific framework of our model, only assuming continuous transition times distributions and the semi-Markov property.
@@
1:subject-specific censoring time $C_{i}\in\bar{\mathbb{R}}^{+}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i,\mathrm{cond}}$, defaults to $t_{i,\mathrm{cond}}=T_{i0}$.
@@
3:Initialize $\ell\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$
@@
5:while $T_{i(\ell-1)}<C_{i}$ and $\{j:(S_{i(\ell-1)},j)\in E\}\neq\emptyset$ do
6:  for each $j$ with $(S_{i(\ell-1)},j)\in E$ do
7:   Draw $T_{i\ell}^{(j)}$ such that $\forall t\geq T_{i(\ell-1)}\vee t_{i,\mathrm{cond}}$
8:      $\log\mathbb{P}(T_{i\ell}^{(j)}>t)=\int_{T_{i(\ell-1)}\vee t_{i,\mathrm{cond}}}^{t}\lambda_{i}^{S_{i(\ell-1)}\to j}(s\mid T_{i(\ell-1)})\,ds$
9:  end for
10:
11:  Set $T_{i\ell}\leftarrow\min_{j}T_{i\ell}^{(j)}$, and $S_{i\ell}\leftarrow\arg\min_{j}T_{i\ell}^{(j)}$
12:  Append $(T_{i\ell},S_{i\ell})$ to $\mathcal{T}_{i}$
13:  $\ell\leftarrow\ell+1$
14:end while
15:
16:if $T_{i(\ell-1)}>C_{i}$ then
17:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$
18:end if
19:return $\mathcal{T}_{i}$
@@
First, consider the case $t_{i,\mathrm{cond}}=T_{i0}$. Let $\ell\geq$ and $t\geq T_{i\ell}$. One has
$\mathbb{P}(T_{i(\ell+1)}>t\mid T_{i\ell},S_{i\ell})$ $=\mathbb{P}\left(\bigcap_{j:(S_{i(\ell-1)},j)\in E}T_{i(\ell+1)}^{(j)}>t\mid T_{i\ell},S_{i\ell}\right),$
$\overset{\mathrel{\perp\!\!\!\perp}}{=}\prod_{j:(S_{i\ell},j)\in E}\mathbb{P}(T_{i(\ell+1)}^{(j)}>t\mid T_{i\ell},S_{i\ell}),$
$=\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i(\ell-1)}\to j}(T_{i\ell},t)\right).$
Furthermore, for any $k^{\prime}\in V$ such that $(S_{i\ell},k^{\prime})\in E$, calling $J_{i\ell}^{(k^{\prime})}\coloneqq\{j:(S_{i\ell},j)\in E\}\setminus\{k^{\prime}\}$, we have
$\mathbb{P}(S_{i(\ell+1)}=k^{\prime}\mid T_{i(\ell+1)}=t,T_{i\ell},S_{i\ell})$
$\propto p\left(T_{i(\ell+1)}^{(k^{\prime})}=t,\forall j\in J_{i\ell}^{(k^{\prime})}\,T_{i(\ell+1)}^{(j)}\geq t\mid T_{i\ell},S_{i\ell}\right)$
$=\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},t)\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},t)\right)$
$\propto\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},t).$
Combining both, we retrieve the same joint density as in Equation 2. If $t_{i,\mathrm{cond}}>T_{i0}$, one can also check that both densities are equal by simply integrating from $t_{\mathrm{cond}}$ instead of $T_{i\ell}$. Note that the condtioning only affects the first transition, as the next time $\forall\ell>0,\,T_{i\ell}\geq t_{i,\mathrm{cond}}\implies\forall\ell>0,\,T_{i\ell}\vee t_{i,\mathrm{cond}}=T_{i\ell}$. ∎
@@ Appendix B Expression of the Semi-Markov likelihood
The proof is short and relies on the fact that, using the semi-Markov property 1
$p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i}\right)$ $=\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell}\right),$
$=\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise, since censoring is assumed non-informative, it can be treated as an independent constant
$\mathbb{P}_{\theta}\left(T_{i(m_{i}+1)}>C_{i}\mid b_{i},X_{i},(T_{im_{i}},S_{im_{i}})\right)=\exp\left(-\sum_{j:(S_{im_{i}},j)\in E}\Lambda_{i,\theta}^{S_{im_{i}}\to j}(T_{im_{i}},C_{i}\mid b_{i},X_{i})\right),$
so we recover the formula cited above, with the convention that an empty sum is 0. ∎
Appendix C Stochastic optimization process on simulated data
Figure 10: Evolution of the parameters during the optimization of the marginal $\log$-likelihood using stochastic gradient ascent. Dotted horizontal lines correspond to the true values, while vertical lines mark the start of each window used to evaluate the stopping criterion.
Appendix D Stochastic optimization process on the PAQUID dataset
Figure 11: Evolution of the parameters during the optimization of the marginal $\log$-likelihood using stochastic gradient ascent. Vertical lines mark the start of each window used to evaluate the stopping criterion.